### v1

@@ 1 Introduction
The evolution from early models based on templates [59, 70] or recurrent neural networks [22, 68] to self-attentive architectures [20, 9, 42] represents significant advancements in image captioning research. These improvements have focused on better connecting visual and textual modalities and incorporating objects and tags at the architectural level [69, 33, 75]. Additionally, there has been a notable emphasis on enhancing the robustness of cross-modal features [34, 5], leading to more accurate captions. Today, image captioning has been integrated into multimodal large language models [31, 37, 11], which demonstrate a strong ability to generate detailed and complex descriptions among other tasks.
@@ 2 Related Work
Image Captioning and Training Stategies. Aligning models with human judgment remains a significant challenge not only in evaluation but also in generation. Early models, ranging from CNN-based encoders and RNNs [64, 22] to the latest fully attentive architectures [20, 42, 9, 34], generate captions by greedily selecting the most probable word from a learned vocabulary. To mitigate error propagation during generation, the beam search algorithm [27] has become widely adopted. This algorithm maintains a set of $k$ most likely sequence candidates and ultimately outputs the most probable sequence from this set.
@@
However, the emergence of pre-trained vision-and-language models like CLIP [45] has highlighted the limitations of traditional metrics for evaluating caption quality and, consequently, for using them as a reward during SCST. In fact, while the use of a CIDEr-based reward can help align generated captions with ground-truth examples, it often reduces the semantic richness of the predicted sentences. To solve this issue, there has been limited exploration of learnable reward models that align references and generated captions without handcrafted metrics [8, 12, 73]. Additionally, there is a growing interest in exploiting the large-scale pre-training of large language models to obtain semantically richer descriptions. In this context, some approaches [47, 40] employ pre-trained language models like GPT-2 and exclusively train specific components, such as cross-attention layers, to capture the complex relationships between images and corresponding textual descriptions. Other solutions [52, 13], instead, directly adapt multimodal large language models to generate more detailed captions.
@@ 3.1 Revisiting CLIP
where $\text{sim}(\cdot)$ is the CLIP-based cosine similarity between visual and textual inputs that are normalized via $\ell_{2}$ normalization, and $\tau$ is a temperature parameter to scale the logits. With the symmetrical loss applied to both image and text encoders, the overall loss function $L_{\mathcal{V},\mathcal{T}}$ is computed as the average of the two.
@@ 3.4 Evaluating Video-Caption Pairs
Given a video $W=\{w^{j}\}_{j=1}^{|W|}$ , where $|W|$ is the number of frames, each fine-grained frame embedding $\mathcal{W}_{f}^{j}$ is obtained as follows:
@@
where $\hat{E_{t}}(\cdot)$ is the CLIP text encoder before the last layer normalization (LN) and linear projection $W$. Each of the $|L|$ token embeddings is used for fine-grained embedding matching, while the [EOS] token serves as the global embedding for coarse-grained embedding matching. Specifically, we define $E_{t}(t)=t_{f}^{\text{{EOS}}}$, which we denote as $t_{c}$ for the sake of notation.
@@ 4.2 Self-Critical Sequence Training with PAC-S++
Given the caption ${\hat{t}}$ and the image $v$, PAC-S++ score is computed and used as the reward $r(\cdot)$ for guiding a policy-gradient reinforcement learning update step:
@@
where $\text{Score}(\cdot)$ is computed as in Eq. 4. Additionally, we consider a variant that takes into account reference captions, thus employing Eq. 5 to compute the reward. To mitigate the variance in the reward signal, a baseline value $b$, computed as the average of the reward of all descriptions generated for $v$, is subtracted from the reward.
@@ 5.1 Implementation Details
[…]
Positive Image-Text Generation. To expand the training dataset with additional positive instances, we use Stable Diffusion [51] for generating new visual data and the BLIP model [32] for generating new textual descriptions. Specifically, to generate images, we employ the model pre-trained on the English image-text pairs of the LAION-5B dataset [55] and fine-tuned at a resolution equal to $512\times 512$ on the LAION-Aesthetics subset [^1] https://laion.ai/blog/laion-aesthetics/, which has been filtered with aesthetic requirements. Throughout the generation process, we utilize a safety checker module to minimize the probability of explicit images. Moreover, we disable the invisible watermarking of the outputs to prevent easy identification of the images as being machine-generated.
Fine-tuning with RL. When assessing the effectiveness of PAC-S++ for fine-tuning captioning models, we employ a standard encoder-decoder Transformer architecture. Specifically, we use three layers in both encoder and decoder, a hidden size of 512, and 8 attention heads. To encode input images, we adopt the CLIP ViT-L/14 visual encoder.
At training stage, we initially pre-train the model with the classical cross-entropy loss for sentence generation. Subsequently, we optimize our model using PAC-S++ in both reference-free and reference-based versions. During cross-entropy pre-training, we train our network with the Adam optimizer [25], a batch size of 1,024, and for up to 20,000 steps. During this phase, we linearly warmup for 1,000 steps, then keep a constant learning rate of $2.5\cdot 10^{-4}$ until 10,000 steps, then sub-linearly decrease until 15,000 steps to $10^{-5}$ and keep the value constant until the end of the training.
@@ 5.2 Evaluating Human Correlation
Kendall $\tau_{b}$ Spearman $\rho$ Kendall $\tau_{b}$ Spearman $\rho$ Kendall $\tau_{b}$ Spearman $\rho$
BLEU-1 [43] - - 12.2 15.9 28.9 37.0
BLEU-4 [43] - - 12.6 16.4 22.4 29.5
ROUGE [35] - - 12.5 16.3 23.8 30.9
METEOR [4] - - 16.4 21.5 27.6 35.7
CIDEr [63] - - 17.3 22.6 27.8 36.1
BERT-S [76] - - 18.2 23.7 29.3 37.8
BERT-S++ [71] - - 15.2 19.8 24.4 31.7
EMScore [58] 23.2 30.3 28.6 37.1 36.8 47.2
28.1 36.4 32.2 41.5 39.8 50.8
PAC-S++ / RefPAC-S++ (+4.9) (+6.1) (+3.6) (+4.4) (+3.0) (+3.6)
Following previous works on these datasets [17], we compute Kendall correlation scores ($\tau_{b}$ and $\tau_{c}$). Results reveal that PAC-S++ outperforms all other metrics in both the reference-free and reference-based setting. Specifically, when comparing our score in a reference-free setting, notable improvements of +3.3 and +2.6 points are observed for $\tau_{c}$ on Flickr8k-Expert and $\tau_{b}$ on Flickr8k-CF, respectively. Similar improvements are evident in the reference-based setting. When comparing PAC-S++ with standard reference-based metrics such as CIDEr or SPICE, the performance gap widens considerably, reaching +11.7/11.8 points with respect to CIDEr on the Flickr8k-Expert dataset.
@@
BLEU-1 [43] 64.6 95.2 91.2 60.7 77.9
BLEU-4 [43] 60.3 93.1 85.7 57.0 74.0
ROUGE [35] 63.9 95.0 92.3 60.9 78.0
METEOR [4] 66.0 97.7 94.0 66.6 81.1
CIDEr [63] 66.5 97.9 90.7 65.2 80.1
BERT-S [76] 65.4 96.2 93.3 61.4 79.1
BERT-S++ [71] 65.4 98.1 96.4 60.3 80.1
TIGEr [21] 56.0 99.8 92.8 74.2 80.7
ViLBERTScore [29] 49.9 99.6 93.1 75.8 79.6
FAIEr [66] 59.7 99.9 92.7 73.4 81.4
MID [24] 67.0 99.7 97.4 76.8 85.2
CLIP-S [17] 55.9 99.3 96.5 72.0 80.9
@@ 5.3 Caption Pairwise Ranking
Accuracy (1 Ref) Accuracy (4 Refs) Accuracy
BLEU-1 [43] 65.7 85.4 60.1
BLEU-4 [43] 66.2 87.0 66.1
ROUGE [35] 54.6 70.4 56.7
METEOR [4] 70.1 82.0 72.9
CIDEr [63] 85.7 94.1 77.9
MID [24] 90.5 90.5 -
CLIP-S [17] 87.2 87.2 -
EMScore [58] - - 89.5
@@ 5.4 Sensitivity to Object Hallucination
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Table 5: Human correlation and accuracy on both image and video captioning datasets varying the visual backbone and the rank size. The reported accuracy results for both the VATEX-EVAL and FOIL datasets are obtained from evaluations conducted with a single reference. PAC-S and RefPAC-S refer to the previous version of the metric introduced in [53]. The overall best results for each backbone are in bold.
LoRA Flickr8k-Expert Flickr8k-CF Composite VATEX-EVAL Pascal-50S FOIL ActivityNet-FOIL
Backbone $r$ Kendall $\tau_{c}$ Kendall $\tau_{b}$ Kendall $\tau_{c}$ Kendall $\tau_{b}$ Accuracy Accuracy Accuracy
PAC-S - 54.3 36.0 55.7 25.1 82.4 89.9 90.1
2 54.3 36.9 58.2 27.4 82.2 90.1 91.1
ViT-B/32 4 54.5 37.0 58.3 28.1 82.3 90.2 91.0
8 54.5 36.9 58.3 27.2 82.0 90.0 90.7
PAC-S++ 16 54.3 37.0 58.5 27.8 81.8 90.1 90.9
RefPAC-S - 55.9 37.6 57.3 31.4 84.7 93.7 93.5
2 55.6 38.0 59.1 32.3 84.4 93.5 93.6
ViT-B/32 4 55.7 37.9 59.1 32.2 84.5 93.5 93.4
8 55.6 38.0 59.2 32.1 84.5 93.6 93.5
RefPAC-S++ 16 55.5 37.9 59.3 32.3 84.5 93.6 93.7
ViT-L/14 PAC-S - 55.5 36.8 56.5 28.9 82.2 91.9 91.2
PAC-S++ 4 57.4 38.5 62.0 32.4 82.4 93.6 92.2
ViT-L/14 RefPAC-S - 57.1 37.7 57.2 31.8 85.0 95.3 94.2
RefPAC-S++ 4 57.9 38.8 61.6 33.4 84.7 95.1 94.3
5.5 Ablation Studies
Effect of the Scaling Factor $w$. The scaling factor, denoted by $w$ in Eq. 4, is used to adjust the scale of the final metric. This adjustment is made to enhance the numerical readability without impacting the ranking of the results. Notably, CLIP-S proposes this analysis and sets $w=2.5$. In our case, due to the different score distributions, we use different $w$ values when using different backbones, as shown in Fig. 4. Specifically, all experiments reported in the preceding tables featuring the ViT-B/32 backbone employ $w$ set at $2.5$, while for ViT-L/14, the value of $w$ is set to $3$.
Low-Rank Analysis. All the analyses conducted so far employ the PAC-S++ version with low-rank adaptation (LoRA). In Table 5, we investigate the effect of different ranks (i.e. 2, 4, 8, 16) across the selected datasets. Overall, the CLIP ViT-B/32 backbone exhibits its best performance with a rank of 4, achieving superior results in both reference-free and reference-based settings. Employing this specific rank dimension, we further compute correlations and accuracy on the CLIP ViT-L/14 backbone.
We also compare the results obtained on both backbones and the previous version of our metric (i.e. PAC-S [53]), in which only the last visual and textual projections of the model are fine-tuned. Notably, except for Pascal-50S and FOIL, the version of our metric with LoRA consistently outperforms the original version, regardless of the rank. For example, employing PAC-S++ with the ViT-B/32 backbone fine-tuned with LoRA (with different rank values) yields superior results compared to its counterpart without LoRA, achieving 58.5 on the Composite dataset (+2.8) and 28.1 on the VATEX-EVAL dataset (+3.0). This demonstrates the effectiveness of this strategy, even in the context of video settings.
@@ 5.5 Ablation Studies
When comparing the results using different backbones, we notice that the ViT-L/14 model consistently outperforms the ViT-B/32 backbone. For instance, on Flickr8k-Expert, we achieve correlation scores of 54.5 and 57.4 with a rank equal to 4 on ViT-B/32 and ViT-L/14, respectively. Comparable results are obtained across all datasets, also considering video-based evaluations and a reference-based setting. These results demonstrate the usefulness of more powerful models to evaluate human correlations and accuracy.
@@ 5.6 Comparisons with Advanced Metrics
All the competitors cited so far include standard metrics, like BLEU or CIDEr, as well as learnable ones, such as CLIP-S. However, more recent metrics have been introduced in the literature that are not directly comparable to our proposed evaluation score due to significant differences in their training methodologies or architectural designs. Nevertheless, the comparison with these recent metrics is worth mentioning, and the results are presented in Table 6.

### v2

@@ 1 Introduction
The evolution from early models based on recurrent neural networks [25, 76] to self-attentive architectures [23, 11, 50] represents significant advancements in image captioning research. These improvements have focused on better connecting visual and textual modalities and incorporating objects and tags at the architectural level [39, 81]. Additionally, there has been a notable emphasis on enhancing the robustness of cross-modal features [40, 6], leading to more accurate captions. Today, image captioning has been integrated into multimodal large language models [36, 44, 13, 8], which show a strong ability to generate detailed and complex descriptions among other tasks.
@@ 2 Related Work
Image Captioning and Training Strategies. Aligning models with human judgment remains a significant challenge not only in evaluation but also in generation. Early models, ranging from CNN-based encoders and RNNs [72, 25] to the latest fully attentive architectures [23, 50, 11, 40], generate captions by greedily selecting the most probable word from a learned vocabulary. To mitigate error propagation during generation, the beam search algorithm [30] has become widely adopted. This algorithm maintains a set of $k$ most likely sequence candidates and ultimately outputs the most probable sequence from this set.
@@
However, the emergence of pre-trained vision-and-language models like CLIP [53] has highlighted the limitations of traditional metrics for evaluating caption quality and, consequently, for using them as a reward during SCST. In fact, while the use of a CIDEr-based reward can help align generated captions with ground-truth examples, it often reduces the semantic richness of the predicted sentences. To solve this issue, there has been limited exploration of learnable reward models that align references and generated captions without handcrafted metrics [10, 14]. Additionally, there is a growing interest in exploiting the large-scale pre-training of large language models to obtain semantically richer descriptions. In this context, some approaches [55, 48] employ pre-trained language models like GPT-2 and exclusively train specific components, such as cross-attention layers, to capture the complex relationships between images and corresponding textual descriptions. Other solutions [60, 15], instead, directly adapt multimodal large language models to generate more detailed captions.
@@ 3.1 Revisiting CLIP
[…]
@@ 3.4 Evaluating Video-Caption Pairs
[…]
@@
[…]
@@ 4.2 Self-Critical Sequence Training with PAC-S++
Given the caption $^{\hat{t}}$ and the image $v$, PAC-S++ score is computed and used as the reward $r(\cdot)$ for guiding a policy-gradient reinforcement learning update step:
@@
[…]
@@ 5.1 Implementation Details
Positive-Augmented Contrastive Learning. As commonly used in other CLIP-based evaluation metrics [20, 27, 66], we employ CLIP ViT-B/32 as backbone to encode images or video frames and textual sentences. Moreover, we report some results using CLIP, SigLIP [80] and SigLIP2 [69] with the ViT-L/14 model to demonstrate the generalizability of our approach to more powerful backbones. To refine the visual and textual representations of the model, we fine-tune visual and textual encoders using the methodology outlined in Sec. 3.2, utilizing the COCO dataset [42] that consists of over 120,000 images accompanied by five captions. In particular, we employ the splits introduced by [25], where 5,000 images are used for validation, 5,000 images are used for testing, and the rest for training. During fine-tuning, we freeze the pre-trained model weights and exploit LoRA [22]. The rank of the decomposition $r$ is set to 4, as it performed favorably in our initial experiments. We use AdamW [45] as optimizer with a learning rate equal to $1\cdot 10^{-4}$ and a batch size of 256. The $\lambda_{v}$ and $\lambda_{t}$ values are selected with a grid search, choosing the combination that provides the best average across datasets. Specifically, we set $\lambda_{v}$ to 0.1 and $\lambda_{t}$ to 0.001, and stop the training stage when the validation loss stops decreasing for 1,500 iterations.
Positive Image-Text Generation. To expand the training dataset with additional positive instances, we use Stable Diffusion [59] for generating new visual data and the BLIP model [38] for generating new textual descriptions. Specifically, to generate images, we employ the model pre-trained on the English image-text pairs of the LAION-5B dataset [63] and fine-tuned at a resolution equal to $\times 512$ 512 on the LAION-Aesthetics subset [^1] https://laion.ai/blog/laion-aesthetics/, which has been filtered with aesthetic requirements. Throughout the generation process, we utilize a safety checker module to minimize the probability of explicit images. Moreover, we disable the invisible watermarking of the outputs to prevent easy identification of the images as being machine-generated.
Fine-tuning with RL. When assessing the effectiveness of PAC-S++ for fine-tuning captioning models, we employ a standard encoder-decoder Transformer architecture. Specifically, we use three layers in both encoder and decoder, a hidden size of 512, and 8 attention heads. To encode input images, we adopt the CLIP ViT-L/14 visual encoder. At training stage, we initially pre-train the model with the classical cross-entropy loss for sentence generation. Subsequently, we optimize our model using PAC-S++ in both reference-free and reference-based versions. During cross-entropy pre-training, we train our network with the Adam optimizer [28], a batch size of 1,024, and for up to 20,000 steps. During this phase, we linearly warmup for 1,000 steps, then keep a constant learning rate of $2.5\cdot 10^{-4}$ until 10,000 steps, then sub-linearly decrease until 15,000 steps to $1\cdot 10^{-5}$ and keep the value constant until the end of the training.
@@ 5.2 Evaluating Human Correlation
Following previous works [20], we compute Kendall correlation scores ($\tau_{b}$ and $\tau_{c}$). Results reveal that PAC-S++ outperforms all other metrics in both the reference-free and reference-based setting. Specifically, when comparing our score in a reference-free setting, notable improvements of +3.3 and +2.6 points are observed for $\tau_{c}$ on Flickr8k-Expert and $\tau_{b}$ on Flickr8k-CF, respectively. Similar improvements are evident in the reference-based setting. When comparing PAC-S++ with standard reference-based metrics such as CIDEr or SPICE, the performance gap widens considerably, reaching +11.7/11.8 points with respect to CIDEr on the Flickr8k-Expert dataset.
@@
$\tau_{b}$ $\rho$ $\tau_{b}$ $\rho$ $\tau_{b}$ $\rho$
BLEU-1 - - 12.2 15.9 28.9 37.0
BLEU-4 - - 12.6 16.4 22.4 29.5
ROUGE - - 12.5 16.3 23.8 30.9
METEOR - - 16.4 21.5 27.6 35.7
CIDEr - - 17.3 22.6 27.8 36.1
BERT-S - - 18.2 23.7 29.3 37.8
BERT-S++ - - 15.2 19.8 24.4 31.7
EMScore 23.2 30.3 28.6 37.1 36.8 47.2
PAC-S++ / 28.1 36.4 32.2 41.5 39.8 50.8
RefPAC-S++ (+4.9) (+6.1) (+3.6) (+4.4) (+3.0) (+3.6)
@@
BLEU-1 64.6 95.2 91.2 60.7 77.9
BLEU-4 60.3 93.1 85.7 57.0 74.0
ROUGE 63.9 95.0 92.3 60.9 78.0
METEOR 66.0 97.7 94.0 66.6 81.1
CIDEr 66.5 97.9 90.7 65.2 80.1
BERT-S 65.4 96.2 93.3 61.4 79.1
BERT-S++ 65.4 98.1 96.4 60.3 80.1
TIGEr 56.0 99.8 92.8 74.2 80.7
ViLBERTScore 49.9 99.6 93.1 75.8 79.6
FAIEr 59.7 99.9 92.7 73.4 81.4
MID 67.0 99.7 97.4 76.8 85.2
CLIP-S 55.9 99.3 96.5 72.0 80.9
@@ 5.3 Caption Pairwise Ranking
Accuracy Accuracy Accuracy
(1 Ref) (4 Refs) (1 Ref)
BLEU-1 65.7 85.4 60.1
BLEU-4 66.2 87.0 66.1
ROUGE 54.6 70.4 56.7
METEOR 70.1 82.0 72.9
CIDEr 85.7 94.1 77.9
MID 90.5 90.5 -
CLIP-S 87.2 87.2 -
EMScore - - 89.5
@@ 5.5 Comparisons with Advanced Metrics
All the competitors cited so far include standard metrics, like BLEU or CIDEr, as well as learnable ones, such as CLIP-S. However, more recent metrics have been introduced in the literature that are not directly comparable to our proposed evaluation score due to significant differences in their training strategies or architectural designs. Nevertheless, the comparison with these recent metrics is worth mentioning, and the results are presented in Table 5, where we show PAC-S++ with both CLIP ViT-B/32 and CLIP ViT-L/14 backbones.
@@
Table 6: Evaluation scores of standard and LLM-based captioners on COCO test set.
Reference-based Reference-free
B-4 M C S RefCLIP-S Polos RefPAC-S++ CLIP-S PAC-S++
Standard Captioning Models
Show and Tell [72] 31.4 25.0 97.2 18.1 0.779 0.585 0.752 0.715 0.654
Show, Attend and Tell [76] 33.4 26.2 104.6 19.3 0.790 0.609 0.766 0.727 0.670
Up-Down [4] 36.7 27.9 122.7 21.5 0.804 0.640 0.778 0.740 0.680
AoANet [23] 39.1 29.0 128.9 22.7 0.813 0.660 0.787 0.753 0.693
$\mathcal{M}^{2}$ Transformer [11] 39.1 29.2 131.2 22.6 0.813 0.629 0.791 0.757 0.699
X-Transformer [50] 39.7 29.5 132.8 23.4 0.819 0.668 0.792 0.762 0.701
DLCT [46] 39.8 29.5 133.8 23.0 0.821 0.670 0.795 0.764 0.705
COS-Net [40] 42.0 30.6 141.1 26.6 0.829 0.692 0.803 0.773 0.712
LLM-based Models
BLIP-2 [37] 43.8 31.7 146.0 25.2 0.838 0.716 0.810 0.782 0.719
IDEFICS-9B [32] 36.8 28.3 125.1 22.0 0.840 0.699 0.803 0.788 0.711
LLaVA-1.5-7B [43] 8.1 28.0 69.6 21.2 0.828 0.666 0.794 0.785 0.707
ShareGPT4V [8] 3.6 17.5 0.0 15.0 0.763 0.650 0.745 0.827 0.726
Humans - 24.1 87.6 21.0 0.822 0.654 0.792 0.782 0.710
Figure 4: Qualitative results showing the differences between captions generated by a captioning model trained on COCO captions (i.e., BLIP-2) and LLM-style captions. Quantitative evaluation shows that CIDEr and CLIP-S scores are sensitive to caption length and stylistic variations, while PAC-S++ demonstrates robustness.